Dearman Incorporated’s current share price is $24.82, with 400,000 shares of stock outstanding. Suppose that Dearman announces that it will invest $875,000 cash in a new project that is expected to increase its free cash flow by $460,000 per year for the indefinite future. If Dearman’s weighted average cost of capital is 16 percent, how will this announcement affect its share price?
Answer #1

Answer

Effect on Share Price ; The market share price will increase as the new investment is to generate $460000 per year for indefinite future

The inflow will increase the value of the Dearman Incorporated as Follow'

Inflow per year for indefinite period = $460000

Rate of return = 16%

Value of Future inflows = 460000/16% = $2875000

Net increase in Value = $2875000 -875000 = $2000000

So there will be net increase of $2000000 in Value of the firm which will positively affect the share price and it will increase

Average increase in Share price = $2000000/400000 shares = $5

so new price will be around = $24.82+$5 = $29.82

Conclusion

As is clear from above calculations as new investment will lead to increase in the overall value of the Dearman so it will increase the current share price of Dearman Incorporate

Hit Thumbs up if satisfied

Have any query mention in comment section please

Thank you